2. Two parallel shafts connected by rolling cylinders turn in the same direction 150 rpm and 100 rpm respectively. The smaller cylinder is 16 in. in diameter. How far apart are the axes of the shafts?


1
Expert's answer
2022-03-04T14:46:01-0500

"\frac{NA}{NB}=\frac{DB}{DA}"

"DB=\frac{NA*DA}{NB}"

"=\frac{150*6}{100}=9in."

rb="\frac{db}{2}=\frac{9}{2}=4.5in"

ra="\frac{DA}{2}=\frac{6}{2}=3in."

distance between axes = ra+rb

=4.5+3

=7.5 in.